If a wire of length 4 meter is carrying a current of 2 Ampere,calculate the magnetic field strength at a distance of 3 meter from the wire
1
Expert's answer
2020-12-06T17:16:32-0500

Answer

Magnetic field due to wire is given by

"B=\frac{\mu_0I}{2\pi r}"

"B=\frac{4\pi \times 10^{-7}\times 2}{2\pi \times3}=1.33 Tesla"